OpenStudy (anonymous):

HW Help Please: Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? What is the percent change from pyramid A to pyramid B?

OpenStudy (abmon98):

1/3*pi*(14)^2*6=392*pi, 3136/392pi=2.5 larger than pyramid A

OpenStudy (anonymous):

What formula is that that you used? @Abmon98

OpenStudy (abmon98):

(392pi/3136)*100=39.2%

OpenStudy (abmon98):

the formula i used is 1/3*pi*r^2*h

OpenStudy (anonymous):

Okay well I was doing the calculations and 14^2= 196 and 196*6=1176....Where did you get 3136/392pi=2.5..... @Abmon98

OpenStudy (abmon98):

1/3*pi*(14)^2*6=392pi

OpenStudy (anonymous):

Oh okay..well why didn't you multiply pi in yet

OpenStudy (anonymous):

Cause that would make it 1230.88

OpenStudy (abmon98):

(1230.88/3136)*100=39.2

OpenStudy (anonymous):

OHHH